if 420(tanA+cotA)=841 then what is value of (116cosA - 58sinA)

Grade:10

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
420(tanA+cotA)=841
420(sinA/cosA +cosA/sinA) =841
420/sinAcosA=841
sin2A=840/841
cos2A=sqrt of (1681/841^2)=41/841
cosA=sqrt of (1+cos2A)/2=21/29
sinA=sqrt of (1-cos2A)/2=20/29
116cosA-58sinA=116(21/29)-58(20/29)=44
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty